OpenStudy (anonymous):

Which of the following sets of integers is a Pythagorean triple? A. 3, 5, 6 B. 5, 12, 13 C. 15, 9, 17 D. 6, 8, 12

OpenStudy (unklerhaukus):

they are pythagorean triples iff the square of the largest number is equal to the sum of the squares of the smaller numbers 3, 5, 6 is not a Pythagorean triple because \[3^2+5^2=9+25=34\neq36=6^2\]

OpenStudy (anonymous):

B is a Pythagorean triple